# Cost Functions in Economic Complexity

### v1

@@ Acknowledgments:
We acknowledge valuable discussions with E. Calò, G. De Marzo, D. Mazzilli, and A. Patelli.
@@ Appendix A Derivation of Cost Functions
Here, we provide the analytical details of the computations for both ECI and NHEFC algorithms. We start with the simplest case, namely ECI.
@@ A.1 ECI
This ensures the existence of a potential function in the variables $Z_{i}$ with the property that $-\nabla_{Z_{i}}U=\dot{Z_{i}}$. From this we recover the explicit expression by integrating Eq. (14):
@@
A.2 NHEFC
For NHEFC, after removing the isolated nodes the map equivalent to Eq. (13) is:
@@ A.2 NHEFC
Appendix B NHEFC: Solution in the case N=2
@@
The iterative dynamics given by $x_{n+1}=f(x_{n})$ can be analyzed using a Lamerey diagram. Observing that the function $f(x)$ has a unique fixed point in the interval $[0,+\infty)$, given by
@@
Furthermore, the function $U(\vec{V})$ reaches a stationary point within the domain $[0,+\infty)^{2}$ if and only if $\vec{V}=(x^{*},x^{*})$, as determined by the condition $\nabla U(\vec{V})=0$. Since $U(\vec{V})\to+\infty$ when $\vec{V}$ approaches the boundary of $[0,+\infty)^{2}$, this unique stationary point is necessarily the absolute minimizer of $U$. Consequently, both sequences $U(\vec{V}^{(2n)})$ and $U(\vec{V}^{(2n+1)})$ converge to $\min U$, ensuring that the entire sequence $U(\vec{V}^{(n)})$ also converges to its minimum value.
@@
Appendix C NHEFC: Regular graphs
@@
In particular, if $K=N-1$, the graph is complete (clique). Notice that, independently of the initial condition $\vec{V}^{(0)}\in(0,+\infty)^{N}$, the iterates are uniformly bounded. One easily verifies that
@@
behaves in a manner analogous to the two-node case. In particular, for any initial condition $X^{(0)}>0$, the sequence $X^{(n)}$ converges to the unique fixed point $V^{*}$ of $f_{K}$ in $[0,\infty)$, where
@@
Hence, the condition for a stationary point, $\nabla U(\vec{V})=0$, is equivalent to the fixed point equation (25). Since $U(\vec{V})$ diverges to +$\infty$+ ∞ as $\vec{V}$ approaches the boundary of $(0,+\infty)^{N}$, we conclude that the fixed point $\vec{V}^{*}$ is indeed the unique absolute minimizer of $U$ in $(0,+\infty)^{N}$.
Appendix D NHEFC: local convergence for general graphs

### v2

@@ Acknowledgments:
We acknowledge valuable discussions with E. Calò, G. De Marzo, D. Mazzilli, A. Patelli, and A. Tacchella.
@@ Appendix A Derivation of Cost Functions
Here, we provide the analytical details of the computations for both ECI and EFC algorithms. We start with the simplest case, namely ECI.
@@ A.1 ECI
This ensures the existence of a potential function in the variables $Z_{i}$ with the property that $\nabla_{Z_{i}}U=\dot{Z_{i}}$. From this we recover the explicit expression by integrating Eq. (14):
@@
A.2 EFC
For EFC, after removing the isolated nodes the map equivalent to Eq. (13) is:
@@ A.2 EFC
Appendix B EFC: Solution in the case N=2
@@
The iterative dynamics given by $x_{n+1}=f(x_{n})$ can be analyzed using a Lamerey diagram. Observing that the function $f(x)$ has a unique fixed point in the interval $0,+\infty)$, given by
@@
Furthermore, the function $U(\vec{V})$ reaches a stationary point within the domain $0,+\infty)^{2}$ if and only if $\vec{V}=(x^{*},x^{*})$, as determined by the condition $\nabla U(\vec{V})=0$. Since $U(\vec{V})\to+\infty$ when $\vec{V}$ approaches the boundary of $0,+\infty)^{2}$, this unique stationary point is necessarily the absolute minimizer of $U$. Consequently, both sequences $U(\vec{V}^{(2n)})$ and $U(\vec{V}^{(2n+1)})$ converge to $\min U$, ensuring that the entire sequence $U(\vec{V}^{(n)})$ also converges to its minimum value.
@@
Appendix C EFC: Regular graphs
@@
In particular, if K=N1K=N-1, the graph is complete (clique). Notice that, independently of the initial condition $\vec{V}^{(0)}\in(0,+\infty)^{N}$, the iterates are uniformly bounded. One easily verifies that
@@
behaves in a manner analogous to the two-node case. In particular, for any initial condition $X^{(0)}>0$, the sequence $X^{(n)}$ converges to the unique fixed point $V^{*}$ of $f_{K}$ in $0,\infty)$, where
@@
Hence, the condition for a stationary point, $\nabla U(\vec{V})=0$, is equivalent to the fixed point equation (25). Since $U(\vec{V})$ diverges to $\infty$ as $\vec{V}$ approaches the boundary of $(0,+\infty)^{N}$, we conclude that the fixed point $\vec{V}^{*}$ is indeed the unique absolute minimizer of $U$ in $(0,+\infty)^{N}$.
Appendix D EFC: local convergence for general graphs